QUESTION NO. 1

Given is the BCD representation of X and Y X=1000 1001 0001 Y=1001 0010

a. Convert X and Y into decimal

b. Reproduce the answer into Hexadecimal and compute X-Y using N-1’s compliment

c. Reproduce the final answer in octa

(need answer as soon as possible)
Answer #1

91 X = 100010010001 Y - 10010010 @Binary to Decimal x = (100010010001), - (142!!) + (Ox 210)+ (0 x 29) + (0128) + (x22) +(6x2Binaue to Hexadecimal X-1000 1000 000 8 J (1000 1001 ०००), (891) 16 8 421 Y-1001 ४५ 0010 2 Y - 12 (10010010), (92)X=(891),6 nys complent nelles complect [ (16) 10-13-N NA3 N = (891) 16 (219310 Hexa todecimal >> (163),0-1-(2193)16 = 4096-after (mills) complect x = 46€ =1902) 10 Y = (6D) 16 = (109)10 Hexto octal converion x = (766) 1902)10 Step:ole Hexa to DecimX= (1902)10 = (3556), ou al convesian » Y = (60)16 = (109)10. Hexa to pecimal (6 Dib. = 6X16+0x160 2 (109)10 . te octal. as D